### v1

# Interpretable Visualizations of Data Spaces for Classification Problems

Abstract
How do classification models “see” our data? Based on their success in delineating behaviors, there must be some lens through which it is easy to see the boundary between classes; however, our current set of visualization techniques makes this prospect difficult. In this work, we propose a hybrid supervised-unsupervised technique distinctly suited to visualizing the decision boundaries determined by classification problems. This method provides a human-interpretable map that can be analyzed qualitatively and quantitatively, which we demonstrate through visualizing and interpreting a decision boundary for chemical neurotoxicity. While we discuss this method in the context of chemistry-driven problems, its application can be generalized across subfields for “unboxing” the operations of machine-learning classification models.
I Introduction
Over the last many years, machine learning (ML) has transformed the way computational materials science research is being done, for example accelerating materials discovery[1], improving the quality of molecular simulations[2, 3], and enabling reliable predictions of materials properties[4]. However, many traditional ML and deep learning methods function as “black boxes,” providing little insight into the underlying physics and chemistry of materials and materials processes. However, in chemical research, the true value of an ML model often goes beyond its ability to make accurate predictions; it also lies in its ability to reveal how the set of input parameters, often carefully chosen to capture the underlying chemistry, relate to the property of interest. Addressing this need for interpretability has driven the development of approaches specifically tailored to uncover structure-property relationships, a pursuit central to the field of Quantitative Structure-Property Relationship (QSPR) modeling.
Many approach QSPR through human-interpretable visualization via unsupervised machine-learning methods. Principal Component Analysis (PCA) is the most standard method for mapping tasks[5]. PCA involves identifying a linear (or non-linear, in the case of kernel PCA) combination of the features that maximizes the variance (i.e. information) retained. Beyond PCA, there exist a slue of other mapping techniques commonly used to qualitatively understand data spaces – t-distributed Stochastic Neighbor Embedding (t-SNE) [6], Uniform Manifold Approximation and Projection (U-MAP) [7], and Isometric Mapping (ISOMAP)[8] focus on the preservation of the local, global, and geometric data structure, respectively, and Locally Linear Embedding (LLE) [9] and multi-dimensional scaling (MDS)[10, 11] preserve distances between neighboring data points. While all of these methods excel at creating informative visualizations and identifying patterns in the structure of the data, the “P” (property) of “QSPR” is only spuriously accounted for, where any relationship between the visualized dimensions and the property of interest is purely coincidental.
So, how do we determine which characteristics of our input data most influence the decisions made by our model? Feature importance algorithms numerically determine the impact of input features in determining the output. Tree-based methods, such as random forests, are most commonly associated with feature importance analysis, as they provide built-in “feature importance” measures, which are typically computed as the total reduction of either the Gini impurity (for classification) or the squared error (for regression) of each of the features [12]. SHAP (SHapley Additive exPlanations), another popular post-training analysis scheme, draws from game theory and involves assigning each feature a value (i.e. SHAP value) representing its contribution to a model’s prediction by considering all possible combinations of feature subsets [13]. Alternatively, many have measured the importance of different features by measuring the performance of feature subsets, either through iteratively adding or removing features[14]. These techniques can excel because they can identify the most relevant descriptors from a large feature set, but they may not be particularly informative when the features are heavily correlated or correspond to abstract concepts. In addition, the amount of information these methods provide beyond a ranking of the features is limited, and they tend to scale poorly when used with high-dimensional data[15].
Figure 1: Relationship between variables and projectors in principal covariates classification. $\mathbf{Z}$ represents the evidence, which is a quantification of class likelihoods, and it can be a tensor in a multilabel, multiclass classification problem.
Thus, while methods from both unsupervised learning and supervised learning address aspects of the QSPR problem, neither can fully do so on their own. However, there exists a family of methods, known as hybrid learning, that draws elements from both supervised and unsupervised learning to more efficiently reveal structure-property relationships. Some of these methods include principal components regression [16], partial least squares regression [17], cluster-wise regression [18], continuum regression [19], and principal covariates regression (PCovR) [20]. Of these methods, PCovR is particularly valuable for QSPR because it is directly formulated as a multi-objective optimization problem in which the regression and dimensionality reduction tasks are optimized simultaneously. PCovR has broad applications within the fields of climate science [21], economic forecasting [22], psychology [23], bioinformatics [24, 25, 26], and molecular crystallization [27]. Similar to PCA, PCovR can also be used to incorporate nonlinear relationships through similarity kernels[28], and can be used with regularization to create sparse solutions [21, 25]. In addition, PCovR has been used to modify other techniques, such as clusterwise regression [29], farthest point sampling, and CUR decomposition [30].
Despite its versatility and wide applicability, PCovR, along with most other hybrid learning methods, is primarily tailored for regression tasks (i.e. where the output is continuous). However, many important materials problems require us to determine relationships between design principles and discrete properties, as are common in classification problems. For classification, however, the lack of a method that integrates dimensionality reduction and classification performance makes QSPR difficult to perform. We propose a new method for QSPR, which we call principal covariates classification (PCovC), that involves using PCovR as an intermediary for predicting class likelihoods. We demonstrate the power of this approach by interpreting decision boundaries in chemical neurotoxicity data using various toxicity assays.
II Results
II.1 Mathematical Basis for Principal Covariates Classification
In the following, we provide a concise description of the underlying mathematics for Principal Covariates Classification (PCovC). We direct interested readers to Helfrecht et al. [28] for a more thorough discussion, including that of data scaling and incorporating non-linearity. Included in the appendices are discussions of other QSPR techniques, namely SHAP analysis[13], linear discriminant analysis[31], and saliency maps[32].
II.1.1 Notation
In this work we assume that the training data is organized into two objects: a ”feature matrix”, $\mathbf{X}$, that contains a set of input variables for a set of observations, and a “target matrix”, $\mathbf{Y}$, that contains the output for each observation. The feature matrix is formatted such that its rows correspond to the individual samples and that its columns correspond to the features. In other words, $\mathbf{X}$ has shape $n_{\mathbf{samples}}\times n_{\mathbf{features}}$. Likewise, the target matrix contains the output labels for each observation, each of which can take an integer value between 00 and $n_{\mathbf{classes}}-1$. Thus, $\mathbf{Y}$ has shape $n_{\mathbf{{samples}}}\times n_{\mathbf{{labels}}}$. We represent the class labels as integers. We denote the latent-space projection of the input data as $\mathbf{T}$, which has shape $n_{\mathbf{samples}}\times n_{\mathbf{components}}$. Finally, we denote $\mathbf{Z}$ as the evidence tensor, which contains confidence scores for each class. $\mathbf{Z}$ has shape of $n_{\mathbf{samples}}\times n_{\mathbf{classes}}\times n_{\mathbf{labels}}$. We assume that $\mathbf{P_{XZ}}$ is determined by a trained classification model, where, depending on the architecture of this model, $\mathbf{Z}$ is used to determine the class of the data points.
In this context, we wish to learn linear mappings between $\mathbf{X}$, $\mathbf{T}$, and $\mathbf{Z}$, which we can define using projection matrices. In this work, we define $\mathbf{P_{AB}}$ as the projector from space $\mathbf{A}$ to space $\mathbf{B}$. Therefore, we can write:
$\mathbf{T}$ $=\mathbf{XP_{XT}}$ (1)
$\mathbf{Z}$ $=\mathbf{TP_{TZ}}=\mathbf{XP_{XT}P_{TZ}}$ (2)
The relationship between these objects is shown in Fig. 1.
II.1.2 Linear Classification
In a classification problem, the goal is to assign an input to one of $n_{\mathbf{classes}}$ discrete classes. In other words, we aim to draw “decision boundaries” that divide the input space into $n_{\mathbf{classes}}$ regions, with each region corresponding to a certain class. In a linear classification problem, these boundaries are lines, or hyperplanes if considering a higher-dimensional space. There are two main approaches to linear classification: deterministic and probabilistic. Deterministic approaches, such as support vector machines, involve finding a discriminant function and use an activation function to map this function to a specific class label. Probabilistic approaches, such as logistic regression, estimate the probability distributions of the input and output and use these probabilities to make classifications.
More formally, linear classifiers make predictions by establishing a linear projection of the input data:
$\mathbf{Z}=\mathbf{XP_{XZ}}$ (3)
and then using activation function $\phi$ to map the continuous $\mathbf{Z}$ to the set of discrete class labels $\mathbf{Y}\in{0,1,2,...n_{\mathbf{classes}}-1}$.
$\mathbf{Y}=\phi(\mathbf{Z})$
In the case of binary classification, we can determine the class by using a step function to apply a threshold to the elements in $\mathbf{Z}$. For example, in a binary support vector machine, the decision is based on the sign of $\mathbf{Z}$:
$\mathbf{Y}[i]=\begin{cases}1&\mathbf{Z}[i]>0\\ 0&\mathbf{else}\end{cases}$ (4)
II.1.3 Principal Covariates Regression
PCovR[20, 28] is a mapping technique that determines a low-dimensional matrix $\mathbf{T}\equiv\mathbf{XP_{XT}}$ that minimizes the function
$\ell=\underbrace{\alpha||\mathbf{X-XP_{XT}P_{XT}^{T}}||^{\mathbf{2}}}_{\text{loss in reconstructing X}}+(1-\alpha)\underbrace{||\mathbf{Y-XP_{XT}P_{TY}}||^{2}}_{\text{loss in predicting Y}}$ (5)
For linear and kernel ridge regression, our projector from $\mathbf{T}$ ($\mathbf{P_{TY}}$) is an analytically solvable function of $\mathbf{X}$ and $\mathbf{Y}$, and thus Eq. (5) is analytically solvable by performing an eigendecomposition of a modified Gram matrix $\mathbf{\tilde{K}}$.
$\mathbf{\tilde{K}}=\mathbf{\alpha XX^{T}}+(1-\alpha)\mathbf{\hat{Y}\hat{Y}^{T}}$ (6)
or, equivalently, a modified covariance matrix $\mathbf{\tilde{C}}$
$\mathbf{\tilde{C}}=\alpha\mathbf{X}^{T}\mathbf{X}+(1-\alpha)\left(\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\mathbf{X}^{T}\mathbf{\hat{Y}\hat{Y}}^{T}\mathbf{X}\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\right)$ (7)
where $\mathbf{\hat{Y}}=\mathbf{XP}_{XY}$ is the linear regression approximation of $\mathbf{Y}$, used to prevent overfitting. This procedure is easily adapted for methods based on data similarity, i.e., kernel models, as detailed in Helfrecht et al. [28].
II.1.4 Principal Covariates Classification
The PCovR loss function (Eq. 5) should not be directly modified to be used in classification tasks – replacing the regression loss with a classification loss often results in a non-analytical function (when the underlying classification algorithm incorporates activation functions into their loss optimization) or is scientifically non-productive, as many classification tasks, particularly multiclass tasks, use numerical values to denote class, where these numerical values are ostensibly arbitrary. We choose to adapt PCovR for classification tasks by leveraging the evidence as our approximation of $\mathbf{Y}$. This yields a new modified Gram matrix $\mathbf{\tilde{K}}$.
$\mathbf{\tilde{K}}=\mathbf{\alpha XX^{T}}+(1-\alpha)\mathbf{ZZ^{T}}$ (8)
whose eigendecomposition (or that of the modified covariance matrix, analogous to Eq. 7) can be used to obtain a relevant low-dimensional map. Here, $\mathbf{Z}$ can be obtained through extracting the weights from a fitted linear classifier and applying Eq.  3. Similar to PCovR, this can be easily adapted for classification models operating on data similarity through kernel methods.
II.2 Exposing decision boundaries in toxicity classifications
Figure 2: Effect of mixing parameter $\alpha$ on the low-dimensional map and classification performance (on an out-of-sample dataset) for the “tox21-ache-p3 (-1)” assay. White points indicate molecules deemed “non-toxic” (Negative) by the assay, blue “toxic” (Positive). Opacity denotes train / test split, with low-opacity points belonging to the training set. In the bottom right of each panel is a confusion matrix showing the accuracy of logistic regression on the resulting map, where “TN” indicates the number of “true negatives”, “FP” indicates the number of “false positives”, and so on. For comparison, logistic regression on the full-dimensional data obtains a confusion matrix of TP=17, TN=505, FP=19, FN=19.
What makes a given molecule toxic? In this case study, we use PCovC to reveal a latent space that discriminates chemicals based on their neurotoxicity. More specifically, we leverage the dataset and classification procedures discussed in Wu et al. [33], which itself contains a processed subset of 68 bioassay endpoints for 7,660 molecules from the Tox21 dataset [34]. For this case study, we use a subset of this dataset, focusing only on the three assays that measure the potential for compounds to disrupt the acetylcholinesterase pathway. Since acetylcholinesterase is responsible for terminating synaptic transmission in the brain, these assays are considered as good indicators of neurotoxicity[35]. Within the dataset, these assays are labeled “tox21-ache-p1 (-1)”, “tox21-ache-p3 (-1)”, and “tox21-ache-p5 (-1)”. We will refer to molecules / data points considered “active antagonists” for a given assay as “toxic” and “inactive” as “non-toxic.”
Figure 3: Molecules from the training set near the decision boundary that demonstrate delineating features of toxicity based on the “tox21-ache-p3 (-1)” assay. The underlying map corresponds to PCovC at $\alpha=0.05$; markers denote different pairs of molecules, with blue and white markers corresponding to toxic and non-toxic molecules. Toxic/non-toxic molecule pairs were chosen to correspond to the molecules closest in PCovC space.
The chemistry and connectivity of these molecules are provided in the form of SMILES strings, from which we generate the RDKit fingerprint [36], which encodes molecular topology by mapping atom connectivity and local structure into a compact, sparse binary representation. We then use PCovC to create a map that delineates chemicals by the “tox21-ache-p3 (-1)” assay.
First, we compare the mapping results of PCovC as a function of our mixing parameter, $\alpha$, in Fig. 2. As expected, a PCA map of this dataset reflects little distinction between toxic and non-toxic molecules, where a subsequent analysis via logistic regression labels all data points as “non-toxic,” i.e., the majority population of the dataset. As $\alpha$ is lowered and the classification task bears more weight in Eq. 5, the separation between classes becomes clearer and our predictive power improves, even superceding that of logistic regression on the full-dimensional data. It is worth noting that the best performance, unlike PCovR, comes from a lower $\alpha$ value than 0.5 – this is likely due to the class imbalance within this dataset, and demonstrates the benefits of the tunability of PCovC.
For comparison, linear discriminant analysis (LDA, App. A.2) produces a map very similar to PCovC at $\alpha=0$, wherein only the labels inform the projector matrix. While this still performs better in subsequent classification tasks than PCA (and marginally worse than analogous PCovC maps), consider the implications of using a single-dimensional map to analyze the decisions made by a classification algorithm. This single-dimensional map can only reflect one dimension of chemical diversity, and all other informative dimensions are lost. Therefore, molecules that are similar in this first dimension, but otherwise dissimilar, cannot be distinguished in subsequent analysis. Furthermore, the orthogonal dimensions, such as the “y” axis in Fig. 2 also carry scientific relevance – these dimensions identify molecular characteristics that do not influence the classification of the data points.
More concretely, consider the analysis of the map in Fig. 3. To identify molecules along the decision boundary, we computed the distances between all “toxic” and “non-toxic” points in the first eight components of PCovC space and found the pairs closest to each other. Shown are eight such pairs, which demonstrate that the model has determined that toxicity can change with small hydrocarbon functionalization, swapping of functional groups, or fluorination. While similar analysis can be achieved through visual inspection (e.g via chemiscope.org[37]), we identified these molecules through simple mathematical operations on the PCovC latent space – enabling higher throughput identification of boundary data points, which can then be used to generate new candidate molecules for toxicity screening.
Figure 4: Maps of the Tox21 dataset considering multiple assays for classification. Here, the leftmost column corresponds to single-assay maps. The center column corresponds to two-assay maps, and the right, center plot corresponds to a three-assay map. Each plot is given for the $\alpha$ that corresponds to the best classification score for the corresponding assays. In the lower right, the PCA map is shown for comparison. LDA is not possible with multitask classification. Marker color corresponds to the values for the three assays, shown in the table in the upper right. Opacity corresponds to the train/test split, with training points shown with transparency.
Yet, drawing definitive conclusions about neurotoxicity from a single assay may be insufficient. By integrating data from multiple assays, we could strengthen conclusions on molecular design and determine confidence intervals. This constitutes a multitask problem that other supervised dimensionality reduction algorithms like LDA are unable to handle. We denote the “tox21-ache-p1 (-1)” assay as “Assay 1”, the “tox21-ache-p3 (-1)” assay (previously discussed) as “Assay 2” and the “tox21-ache-p5 (-1)” assay as “Assay 3”. We can generate multitask PCovC maps, seen in Fig. 4, to incorporate information from each of these assays into a single map.
Of the three assays, Assay 1 is the most sensitive indicator of toxicity. In our postprocessed dataset, 488 of the 5,767 (8.46%) compounds are labeled as toxic by this assay, compared to 193 (3.35%) and 146 (2.53%) compounds for the Assays 2 and 3, respectively. This means that there are significantly many more data points considered “toxic” by Assay 1 that are “non-toxic” according to Assays 2 and 3. The PCovC maps for these singular assays and their combinations are shown in Fig. 4. Note the number of red points, indicating molecules only considered “toxic” by Assay 1, and black points, indicating molecules considered toxic by all assays. Maps generated from solely Assays 2 and 3 position these red points with the “non-toxic” white points. The clearest separation occurs in the map built on Assays 1 and 3, which delineates both decision boundaries clearly and discretely. One can imagine doing similar analysis to that of Fig. 3 to identify the molecules along these boundaries to determine the differentiation by the two assays.
III Discussion
In this work, we demonstrated that PCovC can be a transformative tool for visualizing data spaces that involve classification problems. With the acetylcholinesterase assays from the Tox21 dataset, we center our analysis on molecules that neighbor the decision boundary, and extract from our model that minimal changes in molecular structure can delineate toxic molecules from nontoxic molecules. We observed that, compared to PCA and LDA, combining classification performance and information preservation in a single mapping strategy can unveil decision boundaries while still preserving similarity between data points, and can also lead to improvements in classification performance. In the computational chemical analysis, these aspects can lead to furthering our ability to identify and interpret structure-property relationships, however, we anticipate that this method will be impactful in any discipline where understanding feature importance is scientifically meaningful.
IV Methods
IV.1 PCovC
Obtaining a PCovC map requires a few key steps. First, we must obtain $\mathbf{Z}$ for use in Eq. 8 through fitting a linear classifier with $\mathbf{X}$ and $\mathbf{Y}$. The exact choice of linear classification model is flexible; for the analysis in this work, logistic regression with L2 regularization was chosen for simplicity. The projector from the feature space to the latent space can be mathematically derived through following the procedure in Helfrecht et al. [28], but with the evidence matrix $\mathbf{Z}$ used instead of the linear regression approximation $\mathbf{\hat{Y}}$. When $n_{\mathbf{samples}}>n_{\mathbf{features}}$, one should obtain the projectors by performing an eigendecomposition of a modified Gram matrix $\mathbf{\tilde{K}}$ (Eq.  9).
$\mathbf{P}_{\mathbf{XT}}=\left(\alpha\mathbf{X}^{T}+(1-\alpha)\mathbf{W}\mathbf{Z}^{T}\right)\mathbf{U}_{\tilde{\mathbf{K}}}\mathbf{\Lambda}_{\tilde{\mathbf{K}}}^{-\frac{1}{2}}$ (9a)
$\mathbf{P}_{\mathbf{TX}}=\mathbf{\Lambda}_{\tilde{\mathbf{K}}}^{-\frac{1}{2}}\mathbf{U}_{\tilde{\mathbf{K}}}^{T}\mathbf{X}$ (9b)
In Eq. 9, $\mathbf{U}_{\mathbf{\tilde{K}}}$ and $\mathbf{\Lambda}_{\mathbf{\tilde{K}}}$ are obtained through performing an eigendecomposition of $\mathbf{\tilde{K}}$. When $n_{\mathbf{samples}}<n_{\mathbf{features}}$, it is instead faster to use a modified covariance matrix $\mathbf{\tilde{C}}$ (Eq. 10).
$\mathbf{P}_{\mathbf{XT}}=\left(\mathbf{X}^{T}\mathbf{X}\right)^{-\frac{1}{2}}\mathbf{U}_{\mathbf{\tilde{C}}}^{T}\mathbf{\Lambda}_{\mathbf{\tilde{C}}}^{\frac{1}{2}}$ (10a)
$\mathbf{P}_{\mathbf{TX}}=\mathbf{\Lambda}_{\mathbf{\tilde{C}}}^{-\frac{1}{2}}\mathbf{U}_{\mathbf{\tilde{C}}}^{T}\left(\mathbf{X}^{T}\mathbf{X}\right)^{\frac{1}{2}}$ (10b)
Likewise, $\mathbf{U}_{\mathbf{\tilde{C}}}$ and $\mathbf{\Lambda}_{\mathbf{\tilde{C}}}$ are obtained from performing an eigendecomposition of $\mathbf{\tilde{C}}$.
Finally, we require $\mathbf{P_{TZ}}$, a projector that can map the latent space and the class likelihoods. This projector can be found by extracting the weights from a linear fit between $\mathbf{T}$ and $\mathbf{Z}$. To obtain predicted class labels, one can similarly apply an activation function to $\mathbf{TP_{TZ}}$.
Selecting $\alpha=1$ will still result in a standard PCA decomposition; selecting $\alpha=0$, on the other hand, does not have any intuitive analog, unless if using linear regression to directly predict class likelihoods. In that case, we have $\mathbf{Z}=\mathbf{\hat{Y}}$ and $\mathbf{{P_{TZ}}}$ will be the linear regression solution of $\mathbf{T}$ onto the output $\mathbf{Y}$, thus, PCovC at $\alpha=0$ will produce the same result as a linear regression.
IV.2 Tox21 Preprocessing
This dataset, released in Wu et al. [33], contains Tox21 [34] bioassay data for 68 end points and 7,660 compounds, which are provided as SMILES strings. To better enable a visual analysis of the compounds, we separately convert the compounds into ASE [38] objects using RDKit [36]; 10 compounds were dropped from the dataset as they resulted in a “bad conformer id” error from RDKit. We then follow the preprocessing scheme in Wu et al. [33], in which all compounds other than those with an “inactive” result or the most common active results are not considered for model fitting. For the three acetylcholinesterase assays considered in this study, the number of compounds labeled as active antagonists outnumbered those that were labeled as active agonists by a wide margin. Thus, the active antagonists were kept and the active agonists were discarded. We use the same train-test split provided in the dataset itself; for the rigorous analysis of the “tox21-ache-p3 (-1)” assay, the result is a dataset of 6,357 compounds (5,807 train, 550 test). For the multitask PCovC maps, because we discard chemicals that are not inactive or active antagonists for all three assays, the dataset consists of 5,767 compounds (5,277 train, 490 test). As in the original paper, we then convert the SMILES strings of each of the compounds into RDKit fingerprints with default fingerprint size (2,048 bits).
Appendix A Existing Methods for Data Visualization and Explainability
A.1 Linear and Kernel Principal Components Analysis
PCA is a widely-used linear dimensionality reduction technique that aims to retain variance, or information, present in the feature matrix $\mathbf{X}$. More specifically, it aims to find the orthonormal projection $\mathbf{T}=\mathbf{XP_{XT}}$ that minimizes the reconstruction error $l$ of using the low-dimensional projection to reconstruct the original feature matrix:
$\ell_{proj}=||\mathbf{X-TP_{TX}}||^{2}/n_{\mathbf{{samples}}}$ (11)
It is worth noting that $\mathbf{P_{TX}}$ can be found by invoking the requirement that $\mathbf{P_{XT}}$ is orthonormal: $\mathbf{P_{TX}}=\mathbf{P_{XT}}^{T}$. Additionally, a closed-form solution for $\mathbf{P_{XT}}$ exists, which can obtained through performing an eigendecomposition of the covariance matrix. PCA can be very useful for understanding and visualizing what a dataset may look like, however, it does not explicitly retain discriminative information, so the resulting latent space may not be useful if the directions in the data that are responsible for delineating classes are not the same as those that contain the most variance.
A.2 Linear and Quadratic Discriminant Analysis
Linear Discriminant Analysis (LDA) is a well-known classifier that can also perform supervised dimensionality reduction[31]. Mathematically, LDA involves modeling the density of each class $k$ as a multivariate Gaussian:
$f_{k}(x)=\frac{1}{(2\pi)^{n_{\mathbf{{features}}}/2}|\Sigma_{k}|^{1/2}}e^{-\frac{1}{2}(x-\mu_{k})^{T}\Sigma_{k}^{-1}(x-\mu_{k})}$ (12)
For linear discriminant analysis, we assume that classes have the same covariance matrix $\Sigma_{k}=\Sigma$, which leads to the following linear discriminant function:
$\log P(y=k|x)=x^{T}\Sigma^{-1}\mu_{k}-\frac{1}{2}\mu_{k}^{T}\Sigma^{-1}\mu_{k}+\log P(y=k)$ (13)
For dimensionality reduction, LDA will identify a linear projection that maximizes the separation between classes. One can consider the $n_{\mathbf{classes}}$ class means as vectors in $\mathcal{R}^{n_{\mathbf{features}}}$. There exists some subspace $\mathbf{H}$ that contains each of these vectors that is of dimension at most $n_{\mathbf{classes}}-1$. Directions in the data orthogonal to $\mathbf{H}$ contribute to each class equally and thus do not contribute in distinguishing classes. Thus, computing Euclidean distances in $\mathbf{H}$ will allow us to reduce the data to dimension $n_{\mathbf{classes}}-1$ without sacrificing any information useful for classification. To obtain a lower-dimensional latent space optimal for classification, one can project to a linear subspace $\mathbf{H_{L}}$ that maximizes the variance of the class means after projection. $\mathbf{H_{L}}$ can be obtained through performing PCA on the class means. LDA dimensionality reduction can be useful because it will preserve the inter-class variance in the data, which is useful in discriminating between classes. However, in doing so, the intra-class variance is lost. Thus, understanding relationships between points in the same class can be difficult.
A.3 SHAP Values and Saliency Maps
SHAP is a framework for interpreting machine learning model predictions that draws from Shapley values in cooperative game theory. SHAP assigns each feature a score (or SHAP score) that quantifies the change in the conditional expectation of a model that would occur by including the feature, while also accounting for all possible permutations of features. The SHAP value for a feature $i$ is given by:
$\mathbf{SHAP}_{i}=\sum_{S\subseteq F\setminus i}\frac{|S|!(n_{\mathbf{features}}-|S|-1)!}{n_{\mathbf{features}}!}\ E[f(z)|z_{S}]-E[f(z\setminus i)|z_{S\setminus i}]$ (14)
In Eq. 14, $S$ represents a subset of the feature set $F$ that does not include $i$, and $z_{S}$ represents a vector of features including $S$. SHAP is useful because it can assign quantitative, robust feature importance values for each feature regardless of model architecture, which can then be used to either interpret model predictions or perform feature selection. However, SHAP scales poorly when $n_{\mathbf{features}}$ becomes large[15], and the SHAP values can become less valuable for identifying structure-property relationships when using features that do not carry physical meaning. [13]
Saliency maps can provide an alternative method for understanding the contribution of features to a model’s prediction. Saliency maps involve visualizing the gradient of each feature with respect to an output score, which can be intuitive for images where a saliency map will indicate pixel importance, but may not be useful when features are not inherently spatial.